# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions

### v2

@@ Abstract
Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy. Traditional impact models, whether process-based, statistical, or machine learning, struggle with spatial generalization due to limited observations and concept drift. Recently proposed vision foundation models trained on satellite imagery demand massive compute and are ill-suited for dynamic land-surface prediction. We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions. StefaLand improves predictions on four tasks and five datasets: streamflow, soil moisture, soil composition, and landslide detection compared to prior state-of-the-art. Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications. The model builds on a masked autoencoder backbone that learns deep joint representations of landscape attributes, with a location-aware architecture fusing static and time-series inputs, attribute-based representations that drastically reduce compute, and residual fine-tuning adapters that enhance transfer. While inspired by prior methods, their alignment with geoscience and integration in one model enables robust performance on dynamic land–surface tasks. StefaLand can be pretrained and finetuned on academic compute yet outperforms state-of-the-art baselines and even fine-tuned vision foundation models. To our knowledge, this is the first geoscience land-surface foundation model that demonstrably improves dynamic land-surface interaction predictions and supports diverse downstream applications.
@@ 1 Introduction
Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021, IPCC, 2021). To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content. Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions (Dorigo et al., 2013a). Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively. Soil composition (sand, silt, clay fractions) governs infiltration capacity and root-zone storage, while slope–soil-vegetation interactions directly influence landslide hazards. Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change. Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
@@
Traditionally, these tasks were undertaken by physics-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest (Li et al., 2015). In recent years, there has been a proliferation of data-driven machine learning (ML) models (Solomatine and Ostfeld, 2008). These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs. However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks. A particularly notable gap is that currently established geoscience foundation models have largely been trained on satellite imagery for landcover-identification tasks (Jakubik et al., 2023; 2025, Schmude et al., 2024), which limits their ability to capture dynamic land–atmosphere interactions. As a result, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
@@
A grand challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data. Satellite data are often coarse in resolution and uncertain compared to in-situ measurements. SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010). However, in-situ data, due to the cost of installing instruments and varying policies on data sharing, is only available in high density in certain regions of some developed nations. For example, streamflow gauge data are abundant in the United States, Europe, Australia and Japan, but remain sparse in Africa, South America, and much of Asia (Global Runoff Data Centre, 2020). Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements. As quantified in many studies (Feng et al., 2023), a deep network trained on data from some regions can face substantial performance degradation when applied in data-scarce regions. This occurs partly because there are not enough sites to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies across regions (concept drift). While such limitations hinder traditional supervised ML models, foundation models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
@@
Related Work: In hydrologic and ecosystem predictions, supervised long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) remain a highly popular architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well(Kratzert et al., 2018). Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023, Liu et al., 2024), with evidence of overfitting on continuous signals (Zeng et al., 2022). Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
@@
Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013, Yang et al., 2023). Such an expert-derived design represents a crude practice of unsupervised learning that indicates the importance of understanding the joint data distribution. However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches. Foundation models offer a promising approach to address these spatial generalization challenges. By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains (Zhang et al., 2024).
Existing geoscience foundation models, e.g., TerraMind (Jakubik et al., 2025) and Prithvi (Hsu et al., 2024) from IBM, and Aurora (Bodnar et al., 2025) from Microsoft, have largely focused on satellite imagery, which may not capture the temporal dynamics and physical processes most relevant to land surface predictions. For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation (Jakubik et al., 2023). It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" (Jakubik et al., 2025). However, many critical variables regarding soil and the subsurface are not directly observable from space. In addition, satellite images are noisy and a large portion of the data is redundant across repeated revisits or irrelevant for hydrologic processes. Therefore, it is unclear whether they are of high relevance to land-surface dynamical prediction tasks.
@@
Our contributions: We present Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction. StefaLand improves predictions on streamflow, soil moisture, soil composition, and landslide susceptibility compared to state-of-the-art supervised learning baselines as well as finetuned satellite-image-trained earth foundation models. Especially, it shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks. StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) incorporates a variety of ground-based measurement data, emphasizes relevant land-surface physical processes, drastically reduces compute requirements while retaining global coverage, making it accessible to researchers with modest resources. Pretraining our model required only about 720 V100 GPU hours (could be shorter with more advanced GPUs). The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge. Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models.
@@ 2 Methods
2.1 StefaLand Transformer Architecture
Embedding.
Each dynamic input variable $c$ at time $t$ is independently embedded with a nonlinear two-layer projection:
$z_{t,c}=\text{GELU}(x_{t,c}W_{1,c}+b_{1,c})W_{2,c}+b_{2,c},$ (1)
where $W_{1,c}\in\mathbb{R}^{1\times 64}$ and $W_{2,c}\in\mathbb{R}^{64\times 256}$. Summing across all $C$ dynamic variables yields a per-step embedding:
$z_{t}=\sum_{c=1}^{C}z_{t,c}.$ (2)
Static attributes $s_{i}$ are embedded similarly:
$z_{\text{static},i}=\text{GELU}(s_{i}W_{1,i}+b_{1,i})W_{2,i}+b_{2,i},$ (3)
and concatenated as a special static token, producing the joint sequence:
$Z=[z_{1};z_{2};\ldots;z_{T};z_{\text{static}}].$ (4)
Cross-Variable Group Masking (CVGM).
Let $G=\{g_{1},\ldots,g_{K}\}$ be groups of related variables. During pretraining, a temporal window $\tau,\tau+\ell)$ is sampled, and entire groups are stochastically masked:
$m_{k}\sim\text{Bernoulli}(p_{\text{mask}}).$ (5)
For all $c\in g_{k}$ with $m_{k}=1$, the embedding is replaced by a learned mask vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$. This forces reconstructions to rely on cross-variable dependencies rather than treating drivers independently.
Transformer encoder.
After adding learnable positional encodings $P$, we obtain:
$\tilde{Z}=Z+P.$ (6)
This sequence is processed by $N$ Transformer blocks:
$A^{(\ell)}$ $=\text{MHA}(H^{(\ell-1)}),$ (7)
$\tilde{H}^{(\ell)}$ $=\text{LayerNorm}(H^{(\ell-1)}+A^{(\ell)}),$ (8)
[…]
$H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)}).$ (10)
Decoder and loss.
The final hidden states are passed through a bidirectional LSTM to refine local temporal continuity:
$U=\text{LSTM}(H^{(N)}W_{\text{proj}}+b_{\text{proj}}).$ (11)
Outputs corresponding to masked groups are projected back to the original variable dimensions. The reconstruction loss is
$\mathcal{L}=\sum_{c\in\mathcal{M}}w_{c}\frac{\|\hat{x}_{c}-x_{c}\|_{2}^{2}}{\sigma_{c}^{2}},$ (12)
where $\mathcal{M}$ is the set of masked variables, $\sigma_{c}$ is the standard deviation of variable $c$, and $w_{c}$ is a learnable weight.
@@ 2.2 Pretraining Details
The pretraining dataset is a derived global attribute dataset spanning $\sim$8,000 locations (basins) over 40 years. Variables were chosen to represent the key controls on fluxes of water, energy, momentum, sediment, and nutrients. A complete list of variables, their group assignments, and their sources is provided in Appendix C.
The CVGM is designed so variables with reciprocal or bidirectional causality are masked together, preventing them from acting as predictors for each other. Most groupings are straightforward, such as masking silt and clay percentages together. One exception is soil depth, which we classify as a terrain attribute since it is strongly tied to topographic derivations. By masking and reconstructing at the group level, the model is encouraged to capture cross-domain interactions, such as the coupling between soil texture and climate seasonality. While the grouped-masking strategy has sometimes been employed in multimodality AI models, we did not find it applied in earth foundation models like Terramind, Prithvi, or Aurora, which instead use random dropout, spatial patches, or future periods for masking. Our objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector.
@@ 2.3 Finetuning for Prediction Tasks
Our primary finetuning model, StefaLand-resConn, integrates pretrained embeddings with raw forcings through a residual pathway (Figure D1). Let $E_{t}$ denote the transformer embedding at time $t$, and $x_{t}$ the raw forcings. We compute:
@@
$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o}.$ (15)
Skip connections propagate $E_{t}$ and $r_{t}$ into intermediate layers, allowing general knowledge (from $E_{t}$) to be iteratively refined with task-specific signals ($r_{t}$). This design strengthens spatial generalization while adapting to local variability. To see a visual representation of this D1.
2.4 Foundation model comparisons
For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC (Jakubik et al., 2025, Hsu et al., 2024), using the same fine-tuning heads and training protocols as StefaLand. We coupled them together with our residual connection architecture and finetuned the added units while keeping the foundation models’ weights frozen. Because both TerraMind and PrithviWxC require orders of magnitude larger storage space, data transfer, and fine-tuning costs than StefaLand, we chose comparison tasks selectively, focusing on cases where their pretraining data was potentially most relevant. Accordingly, we compared both models on soil moisture, while streamflow was included only in a more exploratory capacity (only TerraMind). For a rough comparison, StefaLand, TerraMind and PrithviWxC used respectively $\approx 2$, $\approx 11$ and $\approx 27$ TB data during pretraining. Due to the intensive data requirements for PrithviWxC, only surface-level variables likely relevant to land surface interactions, along with all 14 static variables, were used. The full atmospheric variables at differing pressure levels were excluded.

### v3

@@ Abstract
Managing natural resources and mitigating risks from floods, droughts, wildfires, and landslides require models that can accurately predict climate-driven land–surface responses. Traditional models often struggle with spatial generalization because they are trained/calibrated on limited observations and can degrade under concept drift. Recently proposed vision foundation models trained on satellite imagery demand massive compute, and they are not designed for dynamic land surface prediction tasks. We introduce StefaLand, a generative spatiotemporal Earth representation learning model centered on learning cross-domain interactions to suppress overfitting. StefaLand demonstrates especially strong spatial generalization on five datasets across four important tasks: streamflow, soil moisture, soil composition and landslides, compared to previous state-of-the-art methods. The domain-inspired design choices include a location-aware masked autoencoder that fuses static and time-series inputs, an attribute-based rather than image-based representation that drastically reduces compute demands, and residual fine-tuning adapters that strengthen knowledge transfer across tasks. StefaLand can be pretrained and finetuned on commonly-available academic compute resources, yet consistently outperforms state-of-the-art supervised learning baselines, fine-tuned vision foundation models and commercially-available embeddings, highlighting the previously overlooked value of cross-domain interactions and providing assistance to data-poor regions of the world.
Machine Learning, Earth System Science, Hydrology, Land Surface Modeling
@@ 1 Introduction
Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021; IPCC, 2021). To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content. Among these, soil moisture controls ecosystem health and influences land-atmosphere interactions (Dorigo et al., 2013a). Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively. Soil composition (sand, silt, clay fractions) governs infiltration capacity and root-zone storage, while slope–soil-vegetation interactions directly influence landslide hazards. Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change.
@@
Traditionally, these tasks were undertaken by physics-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest (Li et al., 2015). In recent years, there has been a proliferation of data-driven machine learning (ML) models (Solomatine and Ostfeld, 2008). These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs. However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks.
Large vision foundation models have been trained on satellite imagery of earth to facilitate mapping (Tseng et al., 2025), weather prediction (Schmude et al., 2024), air quality (Bodnar et al., 2025) and other geoscientific tasks (Jakubik et al., 2023, 2025). A particularly notable observation is that they have not focused on the interactions among landscape domains (climate, soil, vegetation, terrain, geology). The catchment coevolution hypothesis (Troch et al., 2015) states that terrain, soil, vegetation and climate all coevolve, shaping the landscape we have today. Knowing parts of the landscape domains often allow us to predict the others. This implies that their joint distributions can greatly inform latent processes relevant to the transport of water and materials in the catchments. However, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge, no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
@@
A Grand Challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data. Satellite data are often coarse in resolution and uncertain compared to in-situ measurements. SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010; Liu et al., 2022, 2023b). However, in-situ data, due to the cost of installing instruments and varying policies on data sharing, is only available in high density in certain regions of some developed nations. For example, streamflow gauge data are abundant in the United States, Europe, Australia and Japan, but remain sparse in Africa, South America, and much of Asia (Global Runoff Data Centre, 2020). As quantified in many studies (Feng et al., 2023), a deep network trained on data from some regions can face substantial performance degradation when applied in data-scarce regions. This occurs partly because there are not enough sites to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies across regions (concept drift). While such limitations hinder traditional supervised ML models, foundation or representation learning models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
@@
Related Work: In hydrologic and ecosystem predictions, supervised long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) remain a highly popular architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well (Kratzert et al., 2018). Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023; Liu et al., 2024, 2025b), with evidence of overfitting on continuous signals (Zeng et al., 2022). Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with.
@@
Traditional hydrologic research on ”prediction in ungauged basins” (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013; Yang et al., 2023). Such an expert-derived design represents a crude practice of unsupervised learning that indicates the importance of understanding the joint data distribution. However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches. Representation learning models offer a promising approach to address these spatial generalization challenges. By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains (Zhang et al., 2024).
Recent progress in geoscientific foundation models has been driven by the increasing availability of global Earth system datasets and advances in self-supervised learning, enabling pretrained representations that transfer across tasks and regions (Bommasani et al., 2022; Lacoste et al., 2023). Beyond vision-based approaches, recent work on structured data has shown that foundation models such as TabPFN can learn transferable priors over tabular regression tasks without task-specific training. (Hollmann et al., 2025).
Most existing Earth observation foundation models, including TerraMind (Jakubik et al., 2025), Prithvi (Hsu et al., 2024), Aurora (Bodnar et al., 2025), AlphaEarth (Brown et al., 2025), and Galileo (Tseng et al., 2025), are pretrained primarily on remote sensing imagery and related EO products. These models are effective at capturing surface appearance and spatial patterns, and their representations often act as static or slowly varying landscape descriptors. In particular, AlphaEarth learns global, high-resolution spatial embeddings from large-scale Earth observation data that capture consistent landscape structure across climate and biome gradients, and these embeddings have been shown to transfer effectively as fixed spatial features for a wide range of downstream Earth system prediction tasks, including hydrology. Because AlphaEarth embeddings are globally available and independent of task-specific labels, they provide a strong and widely applicable baseline for evaluating spatial generalization in data-limited settings. However, many variables central to land-surface and hydrologic dynamics, including subsurface properties, storage processes, and long-term temporal interactions, are not directly observable from space and therefore cannot be reliably inferred from imagery alone. Moreover, EO foundation models are typically optimized for surface-level semantic consistency and invariance to transient atmospheric or observational effects, which can further suppress signals that are informative for hydrologic processes, such as moisture-related variability or persistence.
@@
Our contributions: We present the Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), a land-focused geoscientific representation learning model for dynamic land–surface prediction. StefaLand is evaluated on streamflow, soil moisture, soil composition, and landslide susceptibility under spatial holdout regimes. It shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks. StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) incorporates a variety of ground-based measurement data, emphasizes relevant land-surface physical processes, drastically reduces compute requirements while retaining global coverage, making it accessible to researchers with modest resources. Pretraining our model required only about 720 V100 GPU hours (could be shorter with more advanced GPUs). The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-domain interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge. Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models.
Figure 1: Conceptual overview of the StefaLand Structure. Static landscape attributes and dynamic forcings are jointly embedded using a transformer-based masked autoencoder with cross-variable group masking. With relevant dimensionality included.
@@ 2 Methods
2.1 Stefaland Structure
StefaLand is a transformer-based masked autoencoder inspired by bidirectional language models such as BERT (Devlin et al., 2018), designed to jointly embed static landscape attributes and dynamic time-varying forcings within a unified representation. During pretraining, StefaLand randomly masks subsets of the input variables and learns to reconstruct the masked components using the remaining unmasked information. Masking may affect static attributes, temporal variables, or entire variable groups, encouraging the model to learn the joint distribution over heterogeneous land-surface controls rather than relying on any single data source.
A key design choice in StefaLand is cross-variable group masking, in which physically or statistically related variables are masked together rather than independently. This prevents trivial reconstruction through correlated inputs and encourages the model to capture cross-domain linkages, such as interactions between soil texture and climate seasonality or between topography and hydrologic response. Grouped masking strategies have previously been explored in multimodal learning settings, including Presto (Tseng et al., 2023), and here we adapt this idea using a reconstruction loss applied to masked inputs, normalized by variable-wise standard deviations when available. The attention mechanism naturally accommodates missing inputs by suppressing attention weights for masked tokens, allowing StefaLand to flexibly reason over arbitrarily incomplete input configurations. Although more recent masked-model formulations exist, we adopt a BERT-style design for its simplicity, interpretability, and robustness across heterogeneous inputs.
Following pretraining, the encoder produces contextualized embeddings for both static attributes and dynamic forcings, which are reused across downstream tasks through lightweight task-specific heads and residual adaptation pathways, as illustrated in Figure 1.
@@ 2.2 Pretraining Details
The pretraining dataset is a derived global attribute dataset spanning $\sim$8,634 locations (basins) over 40 years. Variables were chosen to represent the key controls on fluxes of water, energy, momentum, sediment, and nutrients. A complete list of variables, their group assignments, and their sources is provided in Appendix  C. A visualization of the pretraining dataset coverage is provided in Appendix C Figure  4.
The cross-variable group masking (CVGM) scheme is designed so that variables with reciprocal or bidirectional relationships are masked together, preventing them from acting as direct predictors for one another. Most groupings are straightforward and common sense, such as masking silt and clay fractions jointly, while a few variables require domain-specific treatment; for example, soil depth is grouped with terrain attributes due to its strong dependence on topographic derivations. The complete set of variables and their group assignments is provided in Table  16. By masking and reconstructing variables at the group level, the model is encouraged to capture cross-domain interactions, such as couplings between soil texture and climate seasonality.
@@ 2.3 Finetuning for Prediction Tasks
We adapt the pretrained StefaLand encoder to downstream prediction tasks using lightweight task-specific heads while keeping the main model weights frozen. This strategy, illustrated in Figure 1, preserves the general-purpose representations learned during pretraining and reduces the risk of overfitting when finetuning on limited task-specific data.
Our primary finetuning configuration, StefaLand-resConn, integrates pretrained embeddings with raw meteorological forcings through a residual adaptation pathway via additive fusion prior to the recurrent decoder. Let $E_{t}$ denote the StefaLand embedding at time $t$, and let $x_{t}$ denote the corresponding raw forcing inputs. We compute:
@@
$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o},$ (3)
where $f_{\text{conv+linear}}(\cdot)$ denotes a shallow convolutional and linear projection block. Residual connections propagate both the pretrained embedding $E_{t}$ and the task-specific signal $r_{t}$ into the recurrent decoder, allowing general spatial knowledge captured during pretraining to be iteratively refined using task-specific temporal information. This design strengthens spatial generalization while retaining flexibility to adapt to local dynamics.
We employ an LSTM decoder for tasks with explicit temporal structure, such as streamflow and soil moisture prediction. For non-temporal or spatially structured tasks, including soil property inference and landslide susceptibility mapping, the temporal decoder is replaced with a task-appropriate adapter head, such as a multilayer perceptron or two-dimensional convolutional neural network (CNN2D). In these cases, frozen StefaLand embeddings are used as contextual features. Across all tasks, only the adapter modules and task heads are updated during finetuning, while the pretrained StefaLand encoder remains frozen, ensuring computational efficiency and stable transfer from the representation learning model.